OpenStudy (anonymous):

Dana is planning to ride her bicycle around the circular race track and wants to know the distance. If the diameter of the track is 2 mile, how far is the distance around the race track?

OpenStudy (4n1m0s1ty):

Its a circular race track, so the distance must be the circumference: C = 2*pi*r where r = diameter/2 Can you solve it now?

Directrix (directrix):

From above ^^ C = 2*pi*r. Twice the radius is the diameter so you could think about the problem as having formula C = pi*d where d = 2 miles . So, the circumference (distance around the track) is ?